# Modeling Language Evolution Using a Spin Glass Approach

### v1

@@ II Modeling Syntactic Evolution
In summary, we assume universal interactions $\{J_{ij\dots k}\}$ as between clusters of syntactic parameters, which are comprised in the most basic model of Implications $\{I_{ij\dots k}\}$ and of random pairwise couplings, or Weights $\{W_{ij}\}$:
@@
Here, $\zeta$ denotes the overall strength of the implications , relative to that of the dyadic interactions. Again, we use the notation $\{J_{ij\dots k}\}$ for consistency with the statistical physics literature [22].
@@
where both $W_{ij}^{S}=W_{ji}^{S}$ and $W_{ij}^{A}=-W_{ji}^{A}$ are taken to be random numbers drawn from the same normal distribution of zero mean and unit variance. The variable $\phi$, ranging as noted above from $\phi=0^{\circ}$ (fully symmetric weights) to $\phi=90^{\circ}$ (fully antisymmetric) is expected to play a crucial role, as symmetry generally tends to promote stability.
In analogy with statistical physics models of glassy Glauber dynamics [23] we update the values of the syntactic parameters in discrete time by choosing at each time step a random sequence (a permutation $i_{p}(i)$ of the $N$ parameters) and using in turn the zero temperature equation
@@
where $S^{\mu}_{j}(t^{\prime})$ represents the state of the $j$-th syntactic parameter in language $\mu$ at time $t$, if $j$ has not yet been updated, $i_{p}(j)>i_{p}(i)$, and at time $t+1$ if already updated; and $\theta_{i}$ is a threshold controlling, in general, the likelihood that parameter $i$ would remain in the unmarked state ($S_{i}=-1$). In our current model, the threshold $\theta_{i}$ and the implication weights $I_{ij\dots k}$ are so set that, in the absence of the random couplings $W_{ij}$, the sum $\sum_{j\dots k}I_{ij\dots k}S^{\mu}_{j}(t^{\prime})\dots S^{\mu}_{k}(t^{\prime})-\theta_{i}$ is exactly zero when the other parameter values are such that parameter $i$ can choose state (hence, with equal probability), and is negative ranging downward to -1, in all other conditions. Since implications are effectively Boolean expressions, this stipulation uniquely determines both $\theta_{i}$ and $I_{ij\dots k}$.
II.1 Symmetric random couplings stabilize syntax
@@
We begin by simulating the future evolution of each of the 58 natural languages in the database, 10 times, with each of 100 distinct random coupling sets, for a total of 1000 runs per language. We collect the number of parameters flipping at each time step and for a first statistical assessment we do not look at each language individually, but rather at averages over a total of 58,000 data points. The results are plotted in Fig. 1.
@@
The black curve in Fig. 1 shows the number of parameters flipping over time, for $\zeta\to\infty$, i.e. subject solely to implicational constraints, those typically emerging in linguistic analyses. Syntactic structures then continue to evolve rapidly, with 20–30 parameters flipping at each time step, which is obvious considering that some of the parameters in the PCM database are not constrained by implications (13 out of 94) and therefore flip with probability 1/2, enabling other parameters to also flip in their wake (see Fig. 10 in the Appendix). In this model, therefore, were it not for the random couplings, natural languages would be fluid, and change over a time scale set, essentially, by the time step.
@@
Introducing random symmetric connections ($\phi=0^{\circ}$) constrains these dynamics: a relatively large number of parameters (e.g., 10–30) still flip during the first or first few time steps, after which the system rapidly gets immobilized, or frozen in a configuration from which it cannot escape (Fig. 1a), regardless of the exact strength of the implications, $\zeta$.
@@
Obviously, freezing reflects the assumption, in this particular framework, that syntactic parameters are influenced solely by endogenous compatibility, in the absence of any external inputs. Consequently, the system rapidly attains a stable equilibrium state; even very weak unknown interactions (corresponding to large $\zeta$) suffice to freeze its evolution. This scenario, albeit a limit case, is interesting insofar as it is not an automatic consequence of the assumptions in the model.
@@
Introducing asymmetry in the random interactions, in fact, can restore fluidity to the syntax network. This is not yet evident with $\phi=30^{\circ}$ (Fig.1b) but it is for higher asymmetry values, like $\phi=60^{\circ}$, Fig.1c, all the way to full anti-symmetry, $\phi=90^{\circ}$, Fig.1d.
@@
Therefore, the way parameters interact with the random couplings has a dramatic effect on the system dynamics: however weak they may be, what is most important is their degree of symmetry.
@@
Partitioning syntactic parameters into classes based on their dependency order — the independent ones in class $D_{0}$, those dependent on a single other parameter in $D_{1}$, those on two, $D_{2}$, on three, $D_{3}$, and so forth — allows a better understanding of how the degree of symmetry in the random couplings affects network dynamics. The implications create a hierarchy among parameters and, when they evolve subject to implications alone, those on top of the hierarchy are completely free, while those in stepwise lower classes are progressively more constrained, and only a smaller fraction flips at every time step. Random symmetric couplings, however weak, make all parameters freeze, irrespective of their position in the hierarchy: syntax evolution is blocked, obstructed by the multitude of mutual constraints (at zero temperature, i.e., in the absence of fast noise). When the random couplings are anti-symmetric, SM Section I and Fig. S1 show that the original hierarchy is expressed again, in a quantitatively similar fashion, at least close to the deterministic limit ($\zeta$ large). Again, the fluidity of the system relies heavily on that of the 13 central $D_{0}$ parameters, and lower classes are progressively more constrained, even though the random couplings are not informed by the class structure.
@@
Figure 3: Phase diagram for the basic model. The heatmap shows the mean number of parameters flipping at the $100^{th}$ time step, on a logarithmic color scale. The angle $\phi$ (in degrees) is on the x-axis, and the strength of the implications $\zeta$ is (logarithmically) on the y-axis. Towards the bottom of the diagram (small $\zeta$), implications become irrelevant, and the model behaves like a simple unstructured disordered system. For nearly symmetric interactions (up to $\phi\approx 30^{\circ}$, marked by the white boundary), dynamics come to halt, as in a spin glass, while more asymmetric interactions lead to fluid behavior. At the top of the diagram, large $\zeta$ and weak but sufficiently symmetric random interactions still stabilize parameters in metastable states of indefinite duration. By increasing asymmetry, the system transitions to a fluid state, though dynamics are slower compared to those occurring with weaker implications and similar degrees of asymmetry.
@@ III A phase diagram for syntax change
The fluidity of the system - whether the network of syntactic parameters evolves or not - appears to be relatively independent of the natural language considered, i.e., the starting binary vector of each simulation, but also, largely, of the strength $\zeta$ of the implications. The critical determinant of fluidity is the degree of symmetry $\phi$ of the random interactions. Its role can be understood by approaching the limit $\zeta\to 0$, in which implications become irrelevant. In such a limit, our model becomes an asymmetric couplings variant of the Sherrington-Kirkpatrick model of a spin glass [22].
@@
There have been relatively few studies of models of spin glass systems subject to asymmetric interactions [24, 25, 26], and they differ in details like the statistics of the random couplings, the update rule or the parametrization of the degree of asymmetry. The couplings in these systems, $J_{ij}$, are typically expressed as a combination of symmetric and antisymmetric components using a slightly different notation, $J_{ij}=J_{ij}^{S}+kJ_{ij}^{A},$ ranging from strictly symmetric ($k=0$) to purely antisymmetric $k\to\infty$. The symmetry factor $k$ is in turn related to the commonly used quantity $\eta$, defined as $\eta:=\langle J_{ij}J_{ji}\rangle/\langle J_{ij}^{2}\rangle,$ which leads to the relations, useful to bridge across studies, $\tan\phi=k=\sqrt{(1-\eta)/(1+\eta)}.$ Note that in these 3 alternative formulations, an average lack of either symmetry or anti-symmetry, that is, simple asymmetry, corresponds to $\phi=45^{\circ}$ or $k=1$ or $\eta=0$.
@@
It has been shown with a variety of approaches and in a number of slightly different variants that, as the degree of coupling asymmetry increases, the system crosses over from a glassy to a fluid regime. The transition may be observed by looking at cycle length [24, 25], correlation entropy [26], or relaxation time [24, 25, 26]. With nearly symmetric couplings, for instance, the cycle length remains independent of system size, while the relaxation time scales as a power of $N$. For highly asymmetric couplings, on the other hand, both cycle length and relaxation time grow exponentially with $N$. With parallel rather than sequential updating, Nutzel and Krey have identified a subtle transition point at $\eta=0.5$, where the attractor length changes significantly [24, 25]. For $\eta>0.5$, the system gets trapped in fixed points or cycles of length 2. For $\eta<0.5$ (more asymmetric couplings), the system enters chaotic dynamics, characterized by long cycles. Crisanti et al, on the other hand, have observed that the system becomes increasingly unpredictable once the asymmetry angle exceeds a critical threshold, and have proposed a scaling law for the mean Shannon entropy around $k_{c}\simeq 0.5$ (i.e., $\eta_{c}\simeq 0.6$, or in our terms $\phi_{c}\simeq 26^{\circ}-27^{\circ}$) [26]. They note that it could be a bona fide phase transition or a simple cross-over. Which of these two slightly divergent findings is in line with our own?
@@
Figure 3 shows a schematic phase diagram in the $\zeta$$\phi$ plane, obtained by counting the average number of flipping parameters at the $100^{th}$ time step, i.e., in the asymptotic regime. The sharp change from a metastable to a fluid regime in the bottom part of the diagram is broadly consistent with the putative phase transition seen in asymmetric spin glass models the correspond to the limit $\zeta\to 0$ of our syntax network. As $\zeta$ increases, the crossover becomes smoother, but still remarkably well defined (considering also the small size of our system, $N=94$) and centered around the same value $\phi=30^{\circ}$. Note that along the vertical axis, the implication strength spans several orders of magnitude, on a logarithmic scale, and that a major decrease in fluidity is apparent, in the antisymmetric limit, only when $\zeta$ gets larger than $\approx 20-50$. The factor $\zeta$ modulates the strength of individual implications, but since there are fewer implicational coefficients (a total of 753, 659 multi-unit coefficients $I_{ij\dots k}$ and 94 thresholds $\theta_{i}$) than pairwise couplings ($94\times 93=8742$, in our fully connected network) and they are also weaker, only above the range $\zeta\approx 20-50$ they can significantly constrain the fluid dynamics. Irrespective of $\zeta$, for sufficiently symmetric interactions, up to $\phi=30^{\circ}$, the system exhibits glassy behavior with short-lived dynamics, and syntactic parameters rapidly settle into one of a disordered multitude of metastable states - a glassy phase. This parallels the slow dynamics and seemingly directionless diversification of grammars. The simulations then indicate that even weak yet sufficiently symmetric random interactions can stabilize multiple states. The behavior for large $\zeta$ is markedly different from the apparent asymptotic limit $\zeta\to\infty$, i.e., no random couplings, where syntactic parameters remain fluid, led by the 13 free ones among them. These findings suggest that incorporating disorder as an additional interaction term — alongside the intrinsic implications — plays a critical role in generating slow dynamics and long-term persistence in language evolution.
@@
Figure 4: A Hebbian term is effective only if strong. Heatmap showing again the average number of parameters flipping at the $100^{th}$ time step, with fully antisymmetric random interactions ($\phi=90^{\circ}$) when adding a stabilizing Hebbian term, as described in the text. The x-axis now denotes the factor $\rho$, which regulates the strength of the Hopfield term, and the y-axis again $\zeta$. Weak Hebbian interactions (up to $\rho\approx 1$), do not affect the fluid behavior prevailing with antisymmetric random interactions, as seen in Fig. 3. As the Hebbian interaction strength increases, the network becomes frozen. The cross-over appears rapid, but note the logarithmic $\rho$ scale.
@@
Figure 5: An apparent OR rule for stabilizing syntax. The heatmap with the number of flipping parameters at the $100^{th}$ time step, for small $\zeta=0.01$, in the $\phi-\rho$ plane. Syntactic parameters stabilize (the blue region) if at least one of two conditions hold: sufficiently symmetric random interactions (up to $\phi\approx 30^{\circ}$) or a strong Hebbian term $\rho\gtrsim 1$. As discussed below, the two conditions are not however equivalent.

### v2

@@ II.1 The purely binary model
In summary, in this first version of the model we assume universal interactions $\{J_{ij\dots k}\}$ between clusters of syntactic parameters, which are comprised in the most basic model of Implications $\{I_{ij\dots k}\}$ and of hidden pairwise couplings, or Weights $\{W_{ij}\}$:
@@
where $\zeta$ denotes the overall strength of the explicit, sometimes complex implications, relative to that of the hidden, though simply dyadic interactions. Again, we use the notation $\{J_{ij\dots k}\}$ for this set of universal conditions on parameter setting, for consistency with the statistical physics literature [sherrington1975solvable].
@@
where both $W_{ij}^{S}=W_{ji}^{S}$ and $W_{ij}^{A}=-W_{ji}^{A}$ are taken to be random numbers drawn from the same normal (Gaussian) distribution of zero mean and unit variance. The variable $\phi$, ranging as noted above from $\phi=0^{\circ}$ (fully symmetric weights) to $\phi=90^{\circ}$ (fully antisymmetric) is expected to play a crucial role, as symmetry generally tends to promote stability.
In analogy with statistical physics models of glassy Glauber dynamics [fisher1963lattice] we update the values of the syntactic parameters in discrete time by choosing at each time step a random sequence (a permutation $i_{p}(i)$ of the $N$ parameters) and using in turn the zero temperature equation (cp. Eq.(4) below)
@@
where $S^{\mu}_{j}(t^{\prime})$ represents the state of the $j$-th syntactic parameter in language $\mu$ at time $t$, if $j$ has not yet been updated, $i_{p}(j)>i_{p}(i)$, and at time $t+1$ if already updated; and $\theta_{i}$ is a threshold controlling, in general, the likelihood that parameter $i$ would remain in the unmarked state ($S_{i}=-1$). How should the thresholds $\theta_{i}$ and the implication weights $I_{ij\dots k}$ be determined? In this first model, they are set so that, in the absence of the hidden couplings $W_{ij}$, the sum $\sum_{j\dots k}I_{ij\dots k}S^{\mu}_{j}(t^{\prime})\dots S^{\mu}_{k}(t^{\prime})-\theta_{i}$ is exactly zero when the other parameter values are such that parameter $i$ can choose state ($\pm 1$ with equal probability), and is negative in all other conditions (then forcing the value -1). As can be seen, this stipulation uniquely determines both $\theta_{i}$ and $I_{ij\dots k}$.
In the more refined model of Sect.VI, we have also run simulations with non-zero temperature, in which essentially Eqs.(3) are satisfied only with a certain probability, which decreases from 1 to 1/2 as the temperature increases. In that second version of the model, the implications are implemented as hard constraints, and the thresholds serve a different function, to set the propensity for each parameter to take the marked +1 value. These points will be taken up later.
II.2 Symmetric hidden couplings stabilize syntax
@@
We begin by simulating the future evolution of each of the 58 natural languages in the database, 10 times, with each of 100 distinct hidden coupling sets, for a total of 1000 runs per language. We collect the number of parameters flipping at each time step and for a first statistical assessment we do not look at each language individually, but rather at averages over a total of 58,000 data points. The results are plotted in Fig. 1.
@@
The black curve in Fig. 1 shows the number of parameters flipping over time, for $\zeta\to\infty$, i.e. subject solely to implicational constraints, those typically emerging in linguistic analyses. Syntactic structures then continue to evolve rapidly, with 20–30 parameters flipping at each time step, which is obvious considering that some of the parameters in the PCM database are not constrained by implications (13 out of 94) and therefore flip with probability 1/2, enabling other parameters to also flip in their wake (see Fig. A.2 in the Appendix). In this model, therefore, were it not for the hidden couplings, natural languages would be fluid, and change over a time scale set, essentially, by the time step.
@@
Introducing hidden symmetric connections ($\phi=0^{\circ}$) constrains these dynamics: a relatively large number of parameters (e.g., 10–30) still flip during the first or first few time steps, after which the system rapidly gets immobilized, or frozen in a configuration from which it cannot escape (Fig. 1a), regardless of the exact strength of the implications, $\zeta$.
@@
Obviously, freezing reflects the assumption, in this particular framework, that syntactic parameters are influenced solely by endogenous compatibility, in the absence of any external inputs. Consequently, the system rapidly attains a stable equilibrium state; even very weak unknown interactions (corresponding by complementarity to large $\zeta$) suffice to freeze its evolution. This scenario, albeit a limit case, is interesting insofar as it is not an automatic consequence of the assumptions in the model.
@@
Introducing asymmetry in the hidden interactions, in fact, can restore fluidity to the syntax network. This is not yet evident for $\phi=30^{\circ}$ (Fig.1b) but it is for higher asymmetry values, like $\phi=60^{\circ}$, Fig.1c, all the way to full anti-symmetry, $\phi=90^{\circ}$, Fig.1d.
@@
Therefore, the way parameters interact with the hidden couplings has a dramatic effect on the system dynamics; and however weak they may be, the key factor is their degree of symmetry, not their strength.
@@
Partitioning syntactic parameters into classes based on their dependency order – the independent ones in class $D_{0}$, those dependent on a single other parameter in $D_{1}$, those on two, $D_{2}$, on three, $D_{3}$, and so forth, following the arrows shown in Fig.A.1 – allows a better understanding of how the degree of symmetry in the hidden couplings affects network dynamics. The implications create a hierarchy among parameters and, when they evolve subject to implications alone, those on top of the hierarchy are completely free, while those in stepwise lower classes are progressively more constrained, and only a smaller fraction flips at every time step. Random symmetric couplings, however weak, make all parameters freeze, irrespective of their position in the hierarchy: syntax evolution is blocked, obstructed by the multitude of mutual constraints (at zero temperature, i.e., in the absence of fast noise). When the hidden couplings are anti-symmetric, Fig. A.2 shows that the original hierarchy is expressed again, in a quantitatively similar fashion, at least close to the deterministic limit ($\zeta$ large). Again, the fluidity of the system relies heavily on that of the 13 central $D_{0}$ parameters, and lower classes are progressively more constrained, even though the hidden couplings are not informed by the class structure.
@@
Figure 3: Phase diagram for the basic model. The heatmap shows the mean number of parameters flipping at the $100^{th}$ time step, on a logarithmic color scale. The angle $\phi$ (in degrees) is on the x-axis, and the strength of the implications $\zeta$ is (logarithmically) on the y-axis. Towards the bottom of the diagram (small $\zeta$), implications become irrelevant, and the model behaves like a simple unstructured disordered system. For nearly symmetric interactions (up to $\phi\approx 30^{\circ}$, marked by the white boundary), dynamics come to halt, as in a spin glass, while more asymmetric interactions lead to fluid behavior. At the top of the diagram, large $\zeta$ and weak but sufficiently symmetric hidden interactions still stabilize parameters in metastable states of indefinite duration. By increasing asymmetry, the system transitions to a fluid state, though dynamics are slower compared to those occurring with weaker implications and similar degrees of asymmetry.
@@ III A phase diagram for syntax change
The fluidity of the system - whether the network of syntactic parameters evolves or not - appears to be relatively independent of the natural language considered, i.e., the starting binary vector of each simulation, but also, largely, of the strength $\zeta$ of the implications. The critical determinant of fluidity is the degree of symmetry $\phi$ of the hidden interactions. Its role can be highlighted by approaching the limit $\zeta\to 0$, in which implications become irrelevant. In such a limit, our model becomes an asymmetric couplings variant of the Sherrington-Kirkpatrick model of a spin glass [sherrington1975solvable].
@@
There have been relatively few studies of models of spin glass systems subject to asymmetric interactions [nutzel1991length, nutzel1993subtle, crisanti1993transition], and they differ in details like the statistics of the hidden couplings, the update rule or the parametrization of the degree of asymmetry. The couplings in these systems, $J_{ij}$, are typically expressed as a combination of symmetric and antisymmetric components using a slightly different notation, $J_{ij}=J_{ij}^{S}+kJ_{ij}^{A},$ ranging from strictly symmetric ($k=0$) to purely antisymmetric $k\to\infty$. The symmetry factor $k$ is in turn related to the commonly used quantity $\eta$, defined as $\eta:=\langle J_{ij}J_{ji}\rangle/\langle J_{ij}^{2}\rangle,$ which leads to the relations, useful to bridge across studies, $\tan\phi=k=\sqrt{(1-\eta)/(1+\eta)}.$ What is important is that these 3 formulations are alternative but equivalent: an average lack of either symmetry or anti-symmetry, that is, simple asymmetry, corresponds to $\phi=45^{\circ}$ or $k=1$ or $\eta=0$.
@@
It has been shown with a variety of approaches and in a number of slightly different variants that, as the degree of coupling asymmetry increases, the system crosses over from a glassy to a fluid regime. The transition may be observed by looking at cycle length [nutzel1991length, nutzel1993subtle], correlation entropy [crisanti1993transition], or relaxation time [nutzel1991length, nutzel1993subtle, crisanti1993transition]. With nearly symmetric couplings, for instance, the cycle length remains independent of system size, while the relaxation time scales as a power of $N$. For highly asymmetric couplings, on the other hand, both cycle length and relaxation time grow exponentially with $N$. With parallel rather than sequential updating, Nutzel and Krey have identified a subtle transition point at $\eta=0.5$ ($\phi=30^{\circ}$), where the attractor length changes significantly [nutzel1991length, nutzel1993subtle]. For $\eta>0.5$, the system gets trapped in fixed points or cycles of length 2. For $\eta<0.5$ (more asymmetric couplings), the system enters chaotic dynamics, characterized by long cycles. Crisanti et al, on the other hand, have observed that the system becomes increasingly unpredictable once the asymmetry angle exceeds a critical threshold, and have proposed a scaling law for the mean Shannon entropy around $k_{c}\simeq 0.5$ (i.e., $\eta_{c}\simeq 0.6$, or in our terms $\phi_{c}\simeq 26^{\circ}-27^{\circ}$) [crisanti1993transition]. They note that it could be a bona fide phase transition or a simple cross-over. Which of these two slightly divergent findings is in line with our own?
@@
Figure 3 shows a schematic phase diagram in the $\zeta$$\phi$ plane, obtained by counting the average number of flipping parameters at the $100^{th}$ time step, i.e., in the asymptotic regime. The sharp change from a metastable to a fluid regime in the bottom part of the diagram is broadly consistent with the putative phase transition seen in asymmetric spin glass models the correspond to the limit $\zeta\to 0$ of our syntax network. As $\zeta$ increases, the crossover becomes smoother, but still remarkably well defined (considering also the small size of our system, $N=94$) and centered around the same value $\phi=30^{\circ}$. Note that along the vertical axis, the implication strength spans several orders of magnitude, on a logarithmic scale, and that a major decrease in fluidity is apparent, in the antisymmetric limit, only when $\zeta$ gets larger than $\approx 20-50$. The factor $\zeta$ modulates the strength of individual implications, but since there are fewer implicational coefficients (a total of 753, 659 multi-unit coefficients $I_{ij\dots k}$ and 94 thresholds $\theta_{i}$) than pairwise couplings ($94\times 93=8742$, in our fully connected network) and they are also weaker, only above the range $\zeta\approx 20-50$ they can significantly constrain the fluid dynamics. Irrespective of $\zeta$, for sufficiently symmetric interactions, up to $\phi=30^{\circ}$, the system exhibits glassy behavior with short-lived dynamics, and syntactic parameters rapidly settle into one of a disordered multitude of metastable states - a glassy phase. This parallels the slow dynamics and seemingly directionless diversification of grammars. The simulations then indicate that even weak yet sufficiently symmetric hidden interactions can stabilize multiple states. The behavior for large $\zeta$ is markedly different from the apparent asymptotic limit $\zeta\to\infty$, i.e., no hidden couplings, where syntactic parameters remain fluid, led by the 13 free ones among them. These findings suggest that incorporating disorder as an additional interaction term — alongside the intrinsic implications — plays a critical role in generating slow dynamics and long-term persistence in language evolution.
@@
Figure 4: A Hebbian term is effective only if strong. Heatmap showing again the average number of parameters flipping at the $100^{th}$ time step, with fully antisymmetric hidden interactions ($\phi=90^{\circ}$) when adding a stabilizing Hebbian term, as described in the text. The x-axis now denotes the factor $\rho$, which regulates the strength of the Hopfield term, and the y-axis again $\zeta$. Weak Hebbian interactions (up to $\rho\approx 1$), do not affect the fluid behavior prevailing with antisymmetric hidden interactions, as seen in Fig. 3. As the Hebbian interaction strength increases, the network becomes frozen. The cross-over appears rapid, but note the logarithmic $\rho$ scale.
@@
Figure 5: An apparent OR rule for stabilizing syntax. The heatmap with the number of flipping parameters at the $100^{th}$ time step, for small $\zeta=0.01$, in the $\phi-\rho$ plane. Syntactic parameters stabilize (the blue region) if at least one of two conditions hold: sufficiently symmetric hidden interactions (up to $\phi\approx 30^{\circ}$) or a strong Hebbian term $\rho\gtrsim 1$. As discussed below, the two conditions are not however equivalent.